i have 51 skittle 15 yellow 6 red 13purple orange 17and green 12 i have to write the ratio of the number of skittles to mass of packet (grams)
1
Expert's answer
2012-10-18T08:18:21-0400
yellow: 15/51=29.4%
red 6/51=11.8%

purple 13/51=25.5%

orange 17/51=33.3%

green 12/51=23.5%